# Batched Line Search Strategy for Navigating through Barren Plateaus in Quantum Circuit Training

### v3

@@ 4 Discerning BPs via monitoring the entropy of subsystems
As mentioned in the introduction, a high vale of entanglement (or von Neumann) entropy [63] is often accompanied with the occurrence of BPs in the optimization landscape. The rapid growing of the entanglement entropy server as an indicator of BP formation, making it a valuable monitoring tool during optimization. Inspired by these findings, we also use the concept of entropy monitoring during the optimization process to characterize the path in the parameter landscape passed during the optimization process.
@@
Here $\rho_{A}=\textrm{tr}_{B}\rho$ is the reduced density matrix where $A$ denotes the subset of qubits that are measured and $B$ is the rest of the system. While in a simulation the evaluation of the entropy is straightforward, it’s value can also be inferred experimentally. The algorithm introduced in [64], called classical shadow estimation, efficiently estimates the expectation value not only of the cost function and its gradients but also the second-Rényi entropy of small-sized subsystems. For practical reasons, we will always normalize the second-Rényi entropy by the Page entropy $S_{Page}$, which is the upper bound of the von Neumann entropy of a Haar random state [65]:
@@ 5 Results
In this section, we present the outcomes of our numerical simulations conducted on variational problems, specifically the estimation of the ground state of an XXX Heisenberg Hamiltonian and the SYC model. In both of these models, the ground state satisfies volume law, indicating that the entanglement entropy of any bi-partition of the state scales with the volume, i.e., $(\rho_{A})\sim|A|$ (refer to Ref. [39] for a comprehensive review of these concepts). Additionally, we share numerical results obtained from quantum gate decomposition problem scenarios. Our findings demonstrate that the developed evolutionary algorithm excels in deep quantum circuit synthesis tasks, outperforming traditional optimization methods such as the well-known Broyden–Fletcher–Goldfarb–Shanno (BFGS) [62] or the ADAM [61] optimization methods. These traditional methods prove ineffective in producing meaningful results for the gate decomposition.
@@
We showcase the efficiency of the evolutionary approach on quantum circuit synthesis problems, particularly in scenarios where our previous findings indicated the presence of local minima leading to the termination of most gradient-based optimizer processes. The gate synthesis problem emerged as a promising candidate to investigate the impact of evolutionary selection on the optimization’s success rate while maintaining a low level of complexity in the cost-function evaluation. This aspect of our numerical experiments is crucial, as optimization with evolutionary selection involves a significant amount of redundant calculations that need to be discarded.
@@
Figure 5: VQE simulations implementing the strategy based on the gradient-free batched search direction with a single agent with $L=128$. The training of the quantum circuits was conducted to find the ground state for the $N=10$-qubit XXX Heisenberg model on a $3$-regular random graph depicted in the subset of sub-figure (a) and for a $10$-qubit SYC model. Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy evaluated on the sites 00 and $1$ (labeled by blue in the insets) for the Heisenberg model. Sub-figures (c) and (d) show the same kind of results for the SYC model. The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-21.28$ eV and $S_{2}/S_{Page}=0.667$. The exact ground state of the SYC model is $E_{0}=-3.82$ eV with normalized entropy $S_{2}/S_{Page}=0.980$. For the initialization we used random parameters.
@@
Though, our numerical results revealed that the addressed system is still too small without the requirement of any special treatments to mitigate the BP effects, provided the VQE energy can be obtained with ideal measurement. Notably, our approach (and other gradient based strategies) is capable to solve the optimization problem without the need for external interventions or monitoring of specific quantities: during the calculations the second-Rényi entropy autonomously remained at low values.
@@
For a circuit with $50$ layers, on the other hand, we could not approximate the ground state with such high quality. Neither the energy nor the entropy [plotted in blue in Fig. 5.(a)-(c)] fully converged to the values corresponding to the exact ground state. The overlap integral at the end of the optimization (terminated at $\sim 3\times 10^{6}$ cost function calls) is $M=0.97$. The results obtained for even shallower circuits rapidly become inaccurate. The explanation behind this observation is related to the expressiveness of the circuits. Considering $10$ qubits, a general wave-function can be described by $10242^{10}=$ 1024 complex amplitudes, resulting in $20462\times 2^{10}-2=$ 2046 free parameters (two parameters are constrained since the state is normalized). The $50$-layered VQE circuit has $2700$ free parameters, while the $100$-layered variant can be tuned with $5400$ free parameters. As one can see, the $50$-layered circuits contain enough parameters to encode any quantum state, but the convergence slows down at the end of the optimization. The $100$-layered circuit converges much faster to the solution, as it is likely over-parameterized.
@@
Figure 6: VQE simulations implementing the strategy based on the gradient-free batched search direction with a single agent with $L=128$. The training of the quantum circuits was conducted to find the ground state for the $N=16$-qubit XXX Heisenberg model on a $3$-regular random graph depicted in the subset of subfigure (c). Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy on sites 00 and $1$ (labeled by blue in the insets) for the Heisenberg model evaluated with shallower circuits consisting of $p=50$ and $100$ layers, while . Subfigures (c) and (d) show the same kind of results obtained with deep circuits consisting of $p=500$ and $1000$ layers. The black numbers and dots represent the evolution of the of the overlap integral between the VQE wave function and the exact ground state defined by Eq.(17). The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-32.84$ eV and $S_{2}/S_{Page}=0.693$. The optimization was initialized with parameters equal to 00.
@@
After examining $10$-qubit systems, we raise the stakes and turn our attention to even more challenging problems. Although it is not a significant challenge for quantum computer simulators to execute circuits beyond $20$ qubits, VQE experiments become quite expensive at this scale due to the large number of cost function evaluations. Therefore, in this work, we limit our considerations to problems up to $16$ qubits. The goal of our VQE experiments is to reproduce the ground state of a randomly generated $16$-qubit XXX Heisenberg model (with $k=3$), as visualized in the inset of Fig. 6.(c). Furthermore, to increase the efficiency of the optimization process, we initiate the training of the circuits with parameter values equal to 00. Since the circuits are applied to the initial state $|\mathbf{0}\rangle$ and initially, all single-qubit rotations will be identities with this choice, the two-qubit entangling gates will lose their effect as well. This is because all control qubits are initially in state 00, so the controlled two-qubit gates incorporated into the circuit will act as an identity gate as well. Consequently, the resulting state will have zero initial entanglement entropy, promising a steep initial descent in the VQE energy. Despite this simplification, our method remains autonomous in avoiding BPs on its own, there is still no need for any external intervention during the optimization.
@@
A general $16$-qubit wave function can be characterized by $1310702^{16}-2=$ 131070 parameters. Consequently, circuits with either $p=50$ or $p=100$ layers are expected to provide much less accurate VQE wave functions than in the previously addressed $10$-qubit case, where the number of trainable parameters in the circuit was sufficiently large. In the $16$-qubit case, the $100$-layered circuit ansatz contains only $9000$ free parameters, which is much less than required to approximate the target state. The left panels of Fig.6 show the numerical results of the conducted VQE experiments with circuits consisting of $p=50$ and $p=100$ layers. As expected, it was not possible to precisely approximate the ground state energy in these instances (VQE energy of $-31.1$ eV with the $50$-layered circuit and $-31.6$ eV with the $100$-layered circuit). The $100$-layered circuit gave significantly better results than those achieved with the $50$-layered circuit. However, neither of these circuits could get close to the exact ground state energy of the model. After more than $10^{8}$ evaluations of the cost function, the VQE wave function still showed a significant difference compared to the exact ground state, and the overlap calculated with the wave function generated by the $100$-layered circuit did not exceed the value of $M=0.75$. The black numbers in Fig.6.(a) indicate the evolution of the overlap with the number of cost function evaluations. As mentioned, these findings were not surprising, as the number of trainable parameters in these circuits was only $4500$ and $9000$ for the $50$- and $100$-layered circuits, respectively. The expressiveness of these circuits lags behind the numbers needed for the accurate approximation of a general $16$-qubit wave function. Therefore, we increased the number of layers in our circuit ansatz by an order.
@@
The right panels of Fig.6 show our numerical results obtained with these deep circuits composed of $p=500$ and $p=1000$ layers. As the plotted results indicate, the developed optimization strategy had no difficulties proceeding with the training of these deep circuits. We still executed only a single agent in these numerical experiments, and optimized $128$ parameters in each iteration. We observed that the efficiency of the optimization varies from run to run. Fig.6 shows the most successful executions, while in several experiments, we observed signs of convergence to a local minimum around the energy level of $\sim-29$ eV. We argue that this issue could be resolved by increasing the number of agents in the process due to the evolutionary selection of the most successful agents. However, due to the overwhelming numerical requirements of the performed VQE simulations, we did not try out this approach in these simulations. We revisit this aspect of the algorithm in a later section by addressing quantum gate synthesis problems, where it becomes essential to avoid local minima to retrieve meaningful results. In these experiments, evolutionary selection will play a crucial role in solving the addressed problems.
@@
To address the optimization problem (2), we explored four strategies available within the SQUANDER package. The Powell-variant of the BFGS optimizer [70] and the GD methods incorporate an inexact line search[40] in each iteration, proving to be quite efficient for smaller problems (e.g., in the synthesis of $3-5$ qubit circuits). Though, the BFGS method, as a second order method is loosing the performance at large parameter count, hence we omit this method from our comparison on VQE experiments. The handcrafted implementations of these optimizers also consider the periodicity of the optimization landscape by appropriately bounding the line search during optimization. Additionally, we included the ADAM optimization algorithm in our tests, although we did not implement the adaptive learning rate strategy proposed in [39]. In Fig.7. we see that the examined optimization engines has similar performance in solving the optimization problem, with the best performance attributed to the ADAM method and to the strategy developed by this work. These two optimizers had much bigger steepness in the cost function curve for both zero and random initializations. The dominant part of the optimization happened in the first quarter of the performed cost function evaluations, while further refinement of the results got significantly slowed down. The steep decrement in the cost function at the beginning can be translated into a reduction in the number of required cost function evaluations to achieve reasonable result, possibly make faster the optimization over BPs.
@@
As our newly developd method reminds the basic concepts of derivative-free approaches, we also explore the capabilities of gradient-free algorithms for comparison. (We took the implementations of the explored gradient-free optimization methods from the SciPy library.) However, we don’t consider our method as a gradient-free approach because our method reproduces the whole cross section of the optimization landscape along single parameter directions. The information on the gradient components also can be accessed with no further computation costs.
@@
Powell’s method was able to solve the optimization problem efficiently, but it exhibited abrupt oscillations in the second-Rényi entropy, which may indicate a loss of efficiency on larger problems. Since the addressed optimization problem is relatively small, it’s too early to draw conclusions about the efficiency of the best-performing optimization strategies at scale. We tried to fill up this gap via further numerical experiments performed on larger systems. Our numerical results showed that by increasing the system size the efficiency of most of the optimization engines start to drop radically. We showcase this tendency via numerical experiments performed on the $16$-qubit Heisenberg model depicted in Fig.6 with $p=500$ circuit layers. This model delivers high degree of computational complexity for a well established numerical investigations.
@@
When initializing the optimization with identity operators (i.e., starting with zero rotational parameters), the second-Rényi entropy begins to evolve from a small value, potentially enabling a steeper decrease in the cost function [as shown in Fig. 9.d)]. Still, our comparison reveals that gradient-based optimizers exhibit significantly lower efficiency in solving the $16$-qubit problem compared to the $10$-qubit VQE problem. This inefficiency is likely linked to the second-Rényi entropy, as the ADAM and GD optimizers, which build up small values of the second-Rényi entropy, are the most efficient among the gradient-based optimizers. In contrast, the gradient descent method with inexact line search enters a region of high entropy early in the optimization, making it highly inefficient. In contrast, the optimization method developed in this work maintains its efficiency observed for the $10$-qubit system during the first quarter of the iterations, rapidly approaching the target value of the optimization [as shown by the blue line in Fig. 9.b)]. However, when the second-Rényi entropy builds up higher values, the steepness of the VQE energy significantly reduces, making only a marginal improvement on the VQE energy. Unfortunately, our numerical experiments were unable to fully recover the target value of the exact ground state. As discussed earlier, the used circuit ansatz may not be optimized for the addressed physical system, potentially lacking the necessary expressiveness to capture the ground state. We leave the study of possible impacts of various circuit structures on the quality of the converged solution for future work. Powell’s method also showed outstanding efficiency in decreasing the cost function, it performed almost equally well as our optimization method. Nevertheless, Powell’s method has a significant limitation. Due to its sequential nature, it becomes impractical for large-scale problems, making it less suitable for VQE tasks at scale. Also, we should notice another limiting well pronounced steps (see the green lines in Fig. 9.a) and b)). In Powell’s method the search directions are determined iteratively from the search directions of the previous step and therefore the optimization along the search directions will affect more and more parameters at once as the optimization iterations go on. Due to our reasoning this is a key component in limiting the efficiency of an optimization method to solve the VQE problem at scale. To overcome this issue, we can restart the solver’s inner state and start over the optimization with search direction aligning with single parameter directions as we do in the optimization method developed in this work. The steps in the green curves in Fig. 9.a) and b) indicates the resets of Powell’s optimization method. This approach allows us to lift the limitation and continue the improvement on the cost function. The other two gradient free optimizers used in our $10$-qubit simulations did not performed well there, hence we omitted them from our larger experiments on $16$ qubits.
@@
In Fig.10, we visualize the evolution of the VQE energy as a function of cost function evaluation count. We observed pronounced oscillations in the cost function for larger learning rates, specifically for $\eta=10^{-2}$ and $\eta=10^{-3}$ (blue and red lines, respectively). Interestingly, the curve for $\eta=10^{-3}$ exhibits the larger oscillations. This phenomenon can be attributed to the optimization landscape’s narrow valleys, which are surrounded by flat territories, consistently with the barren plateau theory (i.e. the dominant part of the landscape is expected to be flat, with narrow valley that are hard to find). A larger learning rate is more likely to overstep these valleys, resulting in oscillations with smaller amplitudes. Conversely, a lower learning rate increases the likelihood of entering these valleys, but may also lead to larger steps during parameter updates (due to the increased gradient inside the valley), causing the algorithm to leave the valley. This reasoning explains the large oscillations in the red line. As the learning rate is further decreased, the oscillations are smoothed out, but the optimization performance is also reduced due to smaller updates in the parameter space (orange and brown lines). Our experiments suggest that learning rates between $\sim 10^{-4}$ and $\sim 10^{-3}$ are the most effective choice.
@@
Interestingly, the oscillations observed during optimization also disappear when updating only a smaller subset of the parameters in each optimization iteration. This means that, similarly to our gradient-free multi-parameter search direction approach developed in this work, we randomly select a finite subset of the parameter space in each iteration, evaluate the associated gradient components of the cost function, and apply the parameter update rule on the chosen subset. To illustrate this tendency, Fig.11 shows the evolution of the VQE energy for various batch sizes $L$. (The batch size refers to the number of the randomly chosen components in the parameters vector.) The circuit ansatz with $500$ layers has $45000$ free parameters. As we decrease the batch size from $45000$ to $500$, the oscillations in the cost function gradually disappear. While the high dimensionality of the optimization landscape makes it challenging to develop an exact explanation, we can offer a qualitative reasoning for this observation. By selecting a small subset of parameters to vary in each iteration, the circuit’s expressiveness gets narrowed down, effectively reducing the risk of getting onto a BP when initiated from zero parameters [20]. By optimizing a low expressive circuit in each iteration of the training process one can effectively avoid BPs during the optimization. However, this procedure does not fully solve the BP issue, as optimization over a selected subspace of the parameter space can only reach the minimum corresponding to that particular subspace. If already close to a local minimum (or on a BP), the improvement in the cost function will be small. These two competing aspects predestinates a batch size for which the optimization is the most efficient. Our numerical results in Fig.11 suggest that a batch size around $50$ is the most efficient choice for gradient descent optimization.

### v4

@@ 4 Discerning BPs via monitoring the entropy of subsystems
As mentioned in the introduction, a high value of entanglement (or von Neumann) entropy [63] is often accompanied with the occurrence of BPs in the optimization landscape. The rapid growth of the entanglement entropy serves as an indicator of BP formation, making it a valuable monitoring tool during optimization. Inspired by these findings, we also use the concept of entropy monitoring during the optimization process to characterize the path in the parameter landscape passed during the optimization process.
@@
Here $\rho_{A}=\textrm{tr}_{B}\rho$ is the reduced density matrix where $A$ denotes the subset of qubits that are measured and $B$ is the rest of the system. While in a simulation the evaluation of the entropy is straightforward, itsits value can also be inferred experimentally. The algorithm introduced in [64], called classical shadow estimation, efficiently estimates the expectation value not only of the cost function and its gradients but also the second-Rényi entropy of small-sized subsystems. For practical reasons, we will always normalize the second-Rényi entropy by the Page entropy $S_{Page}$, which is the upper bound of the von Neumann entropy of a Haar random state [65]:
@@ 5 Results
In this section, we present the outcome of our numerical simulations conducted on variational problems, specifically the estimation of the ground state of an XXX Heisenberg Hamiltonian and the SYC model. In both of these models, the ground state satisfies volume law, indicating that the entanglement entropy of any bi-partition of the state scales with the volume, i.e., $(\rho_{A})\sim|A|$ (refer to Ref. [39] for a comprehensive review of these concepts). Additionally, we share numerical results obtained from quantum gate decomposition problem scenarios. Our findings demonstrate that the developed evolutionary algorithm excels in deep quantum circuit synthesis tasks, outperforming traditional optimization methods such as the well-known Broyden–Fletcher–Goldfarb–Shanno (BFGS) [62] or the ADAM [61] optimization methods. These traditional methods prove ineffective in producing meaningful results for gate decomposition.
@@
We showcase the efficiency of the evolutionary approach on quantum circuit synthesis problems, particularly in scenarios where our previous findings indicated the presence of local minima leading to the termination of most gradient-based optimizer processes. The gate synthesis problem emerged as a promising candidate to investigate the impact of evolutionary selection on the optimization success rate while maintaining a low level of complexity in the cost-function evaluation. This aspect of our numerical experiments is crucial, as optimization with evolutionary selection involves a significant amount of redundant calculations that need to be discarded.
@@
Figure 5: VQE simulations implementing the strategy based on the gradient-free batched search direction with a single agent with $L=128$. The training of the quantum circuits was conducted to find the ground state for the $N=10$-qubit XXX Heisenberg model on a 33-regular random graph depicted in the subset of sub-figure (a) and for a $10$-qubit SYC model. Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy evaluated on the sites 0 and 11 (labeled by blue in the insets) for the Heisenberg model. Sub-figures (c) and (d) show the same kind of results for the SYC model. The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-21.28$ eV and $S_{2}/S_{Page}=0.667$. The exact ground state of the SYC model is $E_{0}=-3.82$ eV with normalized entropy $S_{2}/S_{Page}=0.980$. For the initialization we used random parameters.
@@
Though, our numerical results revealed that the addressed system is still too small without the requirement of any special treatments to mitigate the BP effects, provided the VQE energy can be obtained with ideal measurement. Notably, our approach (and other gradient-based strategies) is capable to solve the optimization problem without the need for external interventions or monitoring of specific quantities: during the calculations the second-Rényi entropy autonomously remained at low values.
@@
For a circuit with $50$ layers, on the other hand, we could not approximate the ground state with such high quality. Neither the energy nor the entropy [plotted in blue in Fig. 5.(a)-(c)] fully converged to the values corresponding to the exact ground state. The overlap integral at the end of the optimization (terminated at $\sim 3\times 10^{6}$ cost function calls) is $M=0.97$. The results obtained for even shallower circuits rapidly become inaccurate. The explanation behind this observation is related to the expressiveness of the circuits. Considering $10$ qubits, a general wave-function can be described by $2^{10}=1024$ complex amplitudes, resulting in $2\times 2^{10}-2=2046$ free parameters (two parameters are constrained since the state is normalized). The $50$-layered VQE circuit has $2700$ free parameters, while the $100$-layered variant can be tuned with $5400$ free parameters. As one can see, the $50$-layered circuits contain enough parameters to encode any quantum state, but the convergence slows down at the end of the optimization. The $100$-layered circuit converges much faster to the solution, as it is likely over-parameterized.
@@
Figure 6: VQE simulations implementing the strategy based on the gradient-free batched search direction with a single agent with $L=128$. The training of the quantum circuits was conducted to find the ground state for the $N=16$-qubit XXX Heisenberg model on a 33-regular random graph depicted in the subset of subfigure (c). Panels (a) and (b) show the evolution of the VQE energy and the normalized second-Rényi entropy on sites 0 and 11 (labeled by blue in the insets) for the Heisenberg model evaluated with shallower circuits consisting of $p=50$ and $100$ layers, while . Subfigures (c) and (d) show the same kind of results obtained with deep circuits consisting of $p=500$ and $1000$ layers. The black numbers and dots represent the evolution of the of the overlap integral between the VQE wave function and the exact ground state defined by Eq.(17). The energy and the normalized entropy of the exact ground state for the Heisenberg system are $E_{0}=-32.84$ eV and $S_{2}/S_{Page}=0.693$. The optimization was initialized with parameters equal to 0.
@@
After examining $10$-qubit systems, we raise the stakes and turn our attention to even more challenging problems. Although it is not a significant challenge for quantum computer simulators to execute circuits beyond $20$ qubits, VQE experiments become quite expensive at this scale due to the large number of cost function evaluations. Therefore, in this work, we limit our considerations to problems up to $21$ qubits. At this point the goal of our VQE experiments is to reproduce the ground state of a randomly generated $16$-qubit XXX Heisenberg model (with $k=3$), as visualized in the inset of Fig. 6.(c). Furthermore, to increase the efficiency of the optimization process, we initiate the training of the circuits with parameter values equal to 0. Since the circuits are applied to the initial state $|\mathbf{0}\rangle$ and initially, all single-qubit rotations will be identities with this choice, the two-qubit entangling gates will lose their effect as well. This is because all control qubits are initially in state 0, so the controlled two-qubit gates incorporated into the circuit will act as an identity gate as well. Consequently, the resulting state will have zero initial entanglement entropy, promising a steep initial descent in the VQE energy. During the optimization , our method remains autonomous in avoiding BPs on its own, there is still no need for any external intervention during the optimization.
@@
A general $16$-qubit wave function can be characterized by $2^{16}-2=131070$ parameters. Consequently, circuits with either $p=50$ or $p=100$ layers are expected to provide much less accurate VQE wave functions than in the previously addressed $10$-qubit case, where the number of trainable parameters in the circuit was sufficiently large. In the $16$-qubit case, the $100$-layered circuit ansatz contains only $9000$ free parameters, which is much less than required to approximate the target state. The left panels of Fig.6 show the numerical results of the conducted VQE experiments with circuits consisting of $p=50$ and $p=100$ layers. As expected, it was not possible to precisely approximate the ground state energy in these instances (VQE energy of 31.1-31.1 eV with the $50$-layered circuit and 31.6-31.6 eV with the $100$-layered circuit). The $100$-layered circuit gave significantly better results than those achieved with the $50$-layered circuit. However, neither of these circuits could get close to the exact ground state energy of the model. After more than $10^{8}$ evaluations of the cost function, the VQE wave function still showed a significant difference compared to the exact ground state, and the overlap calculated with the wave function generated by the $100$-layered circuit did not exceed the value of $M=0.75$. The black numbers in Fig.6.(a) indicate the evolution of the overlap with the number of cost function evaluations. As mentioned, these findings were not surprising, as the number of trainable parameters in these circuits was only $4500$ and $9000$ for the $50$- and $100$-layered circuits, respectively. The expressiveness of these circuits lags behind the numbers needed for the accurate approximation of a general $16$-qubit wave function. Therefore, we increased the number of layers in our circuit ansatz by an order.
@@
The right panels of Fig.6 show our numerical results obtained with these deep circuits composed of $p=500$ and $p=1000$ layers. As the plotted results indicate, the developed optimization strategy had no difficulties proceeding with the training of these deep circuits. We still executed only a single agent in these numerical experiments, and optimized $L=128$ parameters in each iteration. We observed that the efficiency of the optimization varies from run to run. Fig.6 shows the most successful executions, while in several experiments, we observed signs of convergence to a local minimum around the energy level of $\sim-29$ eV. We argue that this issue could be resolved by increasing the number of agents in the process due to the evolutionary selection of the most successful agents. However, due to the overwhelming numerical requirements of the performed VQE simulations, we did not try out this approach in these simulations. We revisit this aspect of the algorithm in a later section by addressing quantum gate synthesis problems, where it becomes essential to avoid local minima to retrieve meaningful results. In these experiments, evolutionary selection will play a crucial role in solving the addressed problems.
@@
Impact of qubit count on the optimization performance:
In the previous section, we demonstrated our optimization technique on a $16$-qubit system. The quantum circuit used in that setup was designed with a layer count that yielded a number of tunable parameters comparable to the degrees of freedom of a $16$-qubit quantum state, enabling a close approximation of the target ground state (see the numbers indicating the overlap integrals in Fig.6. However, as the number of qubits increases, maintaining this correspondence would require an exponential increase in circuit depth. This poses a dual challenge: not only does the expressivity requirement grow exponentially, but also the computational costs of quantum circuit emulation, even if the number of layers is kept constant. To partially address this issue and explore the scalability of our optimization method, we present numerical results for VQE problems with qubit counts ranging from $17$ to $21$, using circuits with a fixed depth of $L=500$ layers. Further investigations of our method’s behavior and performance reported in subsequent sections continue to focus on the $16$-qubit regime, enabling us to effectively approximate the ground state of the targeted system.
Figure 7: (a) Optimization processes to solve Eq.(2) on randomly generated $17$, $18$, 1919, 2020 and 2121-qubit Heisenberg models. The entanglement entropy shown in subfigure (b) stayed at a low value for all of the addressed qubit counts. We executed the optimization method developed in this work with $L=32$.
Figure 7 illustrates that our optimization method, when applied to randomly generated Heisenberg systems, consistently avoided BPs throughout the optimization process. This is supported by the low entanglement entropy values observed across all examined cases. After a sharp initial drop in the VQE energy during the first $7\times 10^{5}$ cost function evaluations, the optimization curves begin to level off, suggesting convergence toward a local minimum. This observation is further supported by analyzing the overlap integrals computed with the exact eigenstates of the addressed physical systems: while the overlap integrals form a shallow distribution taking on value between $10^{-6}$ and $10^{-3}$, one or two values calculated with low-energy eigenstates start to emerge from this distribution. Getting trapped in local minima is a common issue of optimizers. Later in this study we will propose an evolutionary algorithm providing an efficient strategy to escape from them via evolutionary selection rules. We will demonstrate the efficiency of this approach on smaller problems formed for quantum circuit compilation.
These results reinforce the assumption that the optimization process tends to converge toward a local rather than a global minimum. However, due to the lengthy runtime of the optimizer, the convergence was intentionally stopped at an early stage. This decision was driven by the primary objective of our numerical experiments: not to find the final solution, but to evaluate the robustness of the optimization process in the presence of BPs in the optimization landscape.
To make the reported experiments feasible within a practical timeframe, we employed hardware acceleration to significantly speed up quantum circuit emulation. Specifically, we developed a state vector simulator implemented on Groq’s data-flow hardware architecture[70], which we then integrated into the optimization framework used in this study. This integration resulted in up to a $40\times$ speedup in quantum circuit emulation compared to using a 32-core AMD EPYC 7542 processor, within the qubit range considered in our numerical experiments. This performance gain stems from the fact that parallel evaluation of quantum circuits at this scale does not fully utilize the capabilities of modern CPUs. Similarly, GPUs offer limited advantage in this regime, as confirmed by benchmark data provided by QuantumAI’s emulation service [71]. Moreover, our optimization method does not naturally support parallelization, particularly during the line search phase, where cost function evaluations are inherently sequential. As a result, GPU resources cannot be efficiently saturated by the workload required for our experiments. In contrast, Groq’s Language Processing Units are particularly well-suited for this task, as they are optimized for efficient moving of data over the chip enabling fast linear transformations on state vector element pairs. This capability allowed us to significantly accelerate the optimization process, enabling us to report on the results presented in Fig. 7.
@@
To address the optimization problem (2), we explored four strategies available within the SQUANDER package. The Powell-variant of the BFGS optimizer [72] and the GD methods incorporate an inexact line search[40] in each iteration, proving to be quite efficient for smaller problems (e.g., in the synthesis of 353-5 qubit circuits). Though, the BFGS method, as a second order method is loosing the performance at large parameter count, hence we omit this method from our comparison on VQE experiments. The handcrafted implementations of these optimizers also consider the periodicity of the optimization landscape by appropriately bounding the line search during optimization. Additionally, we included the ADAM optimization algorithm in our tests, although we did not implement the adaptive learning rate strategy proposed in [39]. In Fig.8. we see that the examined optimization engines have similar performance in solving the optimization problem, with the best performance attributed to the ADAM method and to the strategy developed by this work. These two optimizers had much bigger steepness in the cost function curve for both zero and random initializations. The dominant part of the optimization happened in the first quarter of the performed cost function evaluations, while further refinement of the results got significantly slowed down. The steep decrement in the cost function at the beginning can be translated into a reduction in the number of required cost function evaluations to achieve reasonable result, possibly making faster the optimization over BPs.
@@
As our newly developed method reminds the basic concepts of derivative-free approaches, we also explore the capabilities of gradient-free algorithms for comparison. (We took the implementations of the explored gradient-free optimization methods from the SciPy library.) However, we don’t consider our method as a gradient-free approach because our method reproduces the whole cross section of the optimization landscape along single parameter directions. The information on the gradient components also can be accessed with no further computation costs.
@@
Powell’s method was able to solve the optimization problem efficiently, but it exhibited abrupt oscillations in the second-Rényi entropy, which may indicate a loss of efficiency on larger problems. Since the addressed optimization problem is relatively small, it’s too early to draw conclusions about the efficiency of the best-performing optimization strategies at scale. We tried to fill up this gap via further numerical experiments performed on larger systems. Our numerical results showed that by increasing the system size the efficiency of most of the optimization engines starts to drop radically. We showcase this tendency via numerical experiments performed on the $16$-qubit Heisenberg model depicted in Fig.6 with $p=500$ circuit layers. This model delivers high degree of computational complexity for exhaustive numerical investigations.
@@
When initializing the optimization with identity operators (i.e., starting with zero rotational parameters), the second-Rényi entropy begins to evolve from a small value, potentially enabling a steeper decrease in the cost function [as shown in Fig. 10.d)]. Still, our comparison reveals that gradient-based optimizers exhibit significantly lower efficiency in solving the $16$-qubit problem compared to the $10$-qubit VQE problem. This inefficiency is likely linked to the second-Rényi entropy, as the ADAM and GD optimizers, which build up small values of the second-Rényi entropy, are the most efficient among the gradient-based optimizers. In contrast, the gradient descent method with inexact line search enters a region of high entropy early in the optimization, making it highly inefficient. In contrast, the optimization method developed in this work maintains its efficiency observed for the $10$-qubit system during the first quarter of the iterations, rapidly approaching the target value of the optimization [as shown by the blue line in Fig. 10.b)]. However, when the second-Rényi entropy builds up higher values, the steepness of the VQE energy significantly reduces, making only a marginal improvement on the VQE energy. Unfortunately, our numerical experiments were unable to fully recover the target value of the exact ground state. As discussed earlier, the used circuit ansatz may not be optimized for the addressed physical system, potentially lacking the necessary expressiveness to capture the ground state. We leave the study of possible impacts of various circuit structures on the quality of the converged solution for future work. Powell’s method also showed outstanding efficiency in decreasing the cost function, it performed almost equally well as our optimization method. Nevertheless, Powell’s method has a significant limitation. Due to its sequential nature, it becomes impractical for large-scale problems, making it less suitable for VQE tasks at scale. Also, we should notice another limiting well pronounced steps (see the green lines in Fig. 10.a) and b)). In Powell’s method the search directions are determined iteratively from the search directions of the previous step and therefore the optimization along the search directions will affect more and more parameters at once as the optimization iterations go on. Due to our reasoning this is a key component in limiting the efficiency of an optimization method to solve the VQE problem at scale. To overcome this issue, we can restart the solver’s inner state and start over the optimization with search direction aligning with single parameter directions as we do in the optimization method developed in this work. The steps in the green curves in Fig. 10.a) and b) indicates the resets of Powell’s optimization method. This approach allows us to lift the limitation and continue the improvement on the cost function. The other two gradient free optimizers used in our $10$-qubit simulations did not performe well there, hence we omitted them from our larger experiments on $16$ qubits.
@@
In Fig.11, we visualize the evolution of the VQE energy as a function of cost function evaluation count. We observed pronounced oscillations in the cost function for larger learning rates, specifically for $\eta=10^{-2}$ and $\eta=10^{-3}$ (blue and red lines, respectively). Interestingly, the curve for $\eta=10^{-3}$ exhibits the larger oscillations. This phenomenon can be attributed to the optimization landscape’s narrow valleys, which are surrounded by flat territories, consistently with the BP theory (i.e. the dominant part of the landscape is expected to be flat, with narrow valleys that are hard to find). A larger learning rate is more likely to overstep these valleys, resulting in oscillations with smaller amplitudes. Conversely, a lower learning rate increases the likelihood of entering these valleys, but may also lead to larger steps during parameter updates (due to the increased gradient inside the valley), causing the algorithm to leave the valley. This reasoning explains the large oscillations in the red line. As the learning rate is further decreased, the oscillations are smoothed out, but the optimization performance is also reduced due to smaller updates in the parameter space (orange and brown lines). Our experiments suggest that learning rates between $\sim 10^{-4}$ and $\sim 10^{-3}$ are the most effective choice.
@@
Interestingly, the oscillations observed during optimization also disappear when updating only a smaller subset of the parameters in each optimization iteration. This means that, similarly to the gradient-free multi-parameter search direction approach developed in this work, we randomly select a finite subset of the parameter space in each iteration, evaluate the associated gradient components of the cost function, and apply the parameter update rule on the chosen subset. To illustrate this tendency, Fig.12 shows the evolution of the VQE energy for various batch sizes $L$. (The batch size refers to the number of the randomly chosen components in the parameters vector.) The circuit ansatz with $500$ layers has $45000$ free parameters. As we decrease the batch size from $45000$ to $500$, the oscillations in the cost function gradually disappear. While the high dimensionality of the optimization landscape makes it challenging to develop an exact explanation, we can offer a qualitative reasoning for this observation. By selecting a small subset of parameters to vary in each iteration, the circuit’s expressiveness gets narrowed down, effectively reducing the risk of getting onto a BP when initiated from zero parameters [20]. By optimizing a low expressive circuit in each iteration of the training process one can effectively avoid BPs during the optimization. However, this procedure does not fully solve the BP issue, as optimization over a selected subspace of the parameter space can only reach the minimum corresponding to that particular subspace. If already close to a local minimum (or on a BP), the improvement in the cost function will be small. These two competing aspects predestinates a batch size for which the optimization is the most efficient. Our numerical results in Fig.12 suggest that a batch size around $50$ is the most efficient choice for gradient descent optimization.